OpenStudy (anonymous):

use the elimination method to solve the systedm of equations. 3u+2v=-15 2u+v=-7

OpenStudy (anonymous):

pl help me im lost

OpenStudy (anonymous):

The elimination method involves subtracting an entire equation from another to remove a variable In order to do this the coefficient of one variable must be the opposite of another, so subtract the second equation by 2 \[2(2u+v=7)\]\[4u+2v = +14\]Subtract this from your first equation \[3u+2v = +15\]\[-(4u+2v = +14)\]or (-u = 1). Plug this back into either equation and solve for v.

OpenStudy (anonymous):

One variable must be EQUAL* to another, not its opposite. And you MULTIPLY an equation by 2, not subtract. If you're still confused let me know

OpenStudy (anonymous):

im

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

elimination example